OpenStudy (anonymous):

Simplify: the square root of 2 (x − the square root of 5) 2 times x times the square root of 2 10 the square root of 2 times x the square root of 2 times x − the square root of 10 x times the square root of 2 − the square root of 10 Simplify: the square root of 6 (7 the square root of 3 + 6) 7 the square root of 108 21 the square root of 2 + 6 the square root of 6 7 the square root of 18 + 6 the square root of 6 126 the square root of 2

OpenStudy (anonymous):

\[\Huge\bf \color{yellow}{Welcome~\to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~\to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~\to~\color{purple}{Open}}\color{blue}{Study!!!!}\]

OpenStudy (anonymous):

Is the first question \[\Large 2(x-\sqrt5)\]?

OpenStudy (anonymous):

Sqrt(2) (x-sqrt(5)

OpenStudy (anonymous):

@doulikepiecauseidont

OpenStudy (anonymous):

\[\Large \sqrt{a}*\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

Multiply sqrt(2) with x-sqrt(5))?

OpenStudy (anonymous):

\[\Large \sqrt{2}(x-\sqrt{5})\]

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So it would be sqrt(10x)?

OpenStudy (anonymous):

@doulikepiecauseidont

OpenStudy (anonymous):

It would be \[\Large 2\sqrt{x}-\sqrt{10}\]

OpenStudy (triciaal):

@doulikepiecauseidont no rt2*x = x*rt 2

OpenStudy (triciaal):

@HishamNafiseh the correct answer for qquestion 1 would be the 4th option

OpenStudy (triciaal):

typo question

OpenStudy (triciaal):